# Statistical modeling of SOFR term structure

### v1

@@ 4.5 Incorporation of user’s views
We apply the calibration procedure from Section 4.2 of [27], which allows us to incorporate both short term forecasts as well as long-term views in both the macroeconomic model as well as the forward curve model described in Sections 4.3 and 4.4 above. The calibration is achieved by choosing the sequence of vectors $a_{t}$ in (17) so that the values of $x_{t}$ have desired median values at given points in time. Since both the macro model and the forward curve model are stationary, the model calibration according to Section 4.2 of [27] is particularly easy.
@@
Table 11: Long-term median for the macroeconomic risk factors $y$. The $2.5\%$ for $L$ is user-specified. The $2.00\%$ for $I$ corresponds to the inflation target set by the FOMC. The long-term median for $G$ is the median computed from historical data.
$L$ $I$ $G$
$c$ $2.50$% $2.00$% $2.85$%
@@ 5.1 Simulation of the macroeconomy and the forward curve
Figure 8: Simulated scenarios for $L$ (in decimals), $I$ (in percentage), and $G$ (in percentage). The plots show the monthly historical data with a single simulated scenario, the median, and the 95% confidence band.
@@ 5.2 Simulating SOFR derivatives payouts
Table 13: Payouts in USD for long positions in SOFR derivatives. The payouts follow the CME convention where the quantities in the formulas appear to have flipped signs; see Remark 1. The maturities are $t_{1}$ for the futures, call and put while the maturity of the swaption is denoted by $T$. The strikes are denoted by $X$.
@@
Three-month SOFR futures $[F_{t}^{b}(t_{0},t_{1})-R(t_{0},t_{1})]\times 250000$
@@
Put on three-month SOFR futures $(F_{t_{1}-1}(t_{1},t_{2})-X)^{+}\times 250000$
@@
If exercised, a swaption, with strike $X$ and maturity $T$, converts to a short position in an OIS with swap rate $X$ and $K$ yearly payments; see [15, 16]. Assuming perfectly liquid zero-coupon bonds, the market price of such an OIS would be
@@ 5.3 Indifference pricing of SOFR derivatives
A common approach to derivative pricing is to construct a market model together with a martingale measure which is then used for pricing derivatives by taking expectations of their cashflows. Fixing a pricing measure, however, leads to a linear pricing rule, which contradicts the nonlinearities observed in practice. Such problems are addressed explicitly by the techniques of incomplete markets, where one first develops a term structure model under the “statistical” measure and then defines prices according to the indifference principle; see e.g. [9], [33] and their references. This section uses the model developed in Sections 2,  3, and 4 in indifference pricing of the derivatives described in the previous section. The exercise below is quite simple in that the only “hedging instrument” is the money market account represented by the overnight Repo market. More sophisticated hedging strategies will be built in the follow-up paper.
@@
For the call and put option, where the price is received at the time of pricing, we use the SOFR overnight roll-over account as a numeraire. This permits computing the selling price again analytically without numerical line search. The selling price is the solution $\alpha$ to the equation
$EU\left(\bar{w}+\alpha-\frac{c}{\prod_{t=t_{0}}^{t_{1}-1}(1+r_{t}\delta)}\right)=EU\left(\bar{w}\right).$ (20)
With the exponential utility, we obtain
$\alpha=\frac{1}{\rho}\ln E\exp\left\{\frac{\rho c}{\prod_{t=t_{0}}^{t_{1}-1}(1+r_{t}\delta)}\right\}.$ (21)
The indifference prices $\alpha$ in (19) and (21) can be computed by numerically evaluating the expectations. In the numerical computations below, we will use Monte Carlo with antithetic sampling. That is, we simulate a random sample of $N$ scenarios and obtain another $N$ scenarios by reflecting the underlying Gaussian process.
@@ 6 Arbitrage
The following theorem is from [23, Theorem 3], but one should note that part g of the statement there has a typo: the symbol $Q_{n}$ should be replaced by $\bar{Q}_{n}$. An alternative proof and an extended statement can be found in [34, Remark 2.120]. The ${{\cal F}_{t}}$-conditional support of a random variable $X\in L^{0}(\Omega,{\cal F},P;\mathbbm{R}^{d})$ is the smallest closed-valued ${\cal F}_{t}$-measurable mapping $\mathop{\rm supp}\nolimits_{t}X:\Omega\rightrightarrows\mathbbm{R}^{d}$ such that $X\in\mathop{\rm supp}\nolimits_{t}X$ almost surely. The convex hull and the relative interior of a set $C\subseteq\mathbbm{R}^{d}$ will be denoted by $\mathop{\rm co}C$ and $\mathop{\rm ri}C$, respectively.
@@ 6.1 Arbitrage in zero-coupon bonds
By [34, Theorem 2.43.4], the continuity of $H$ implies
@@
The claim now follows from the fact that $\mathop{\rm ri}\mathop{\rm co}\mathop{\rm cl}C=\mathop{\rm ri}\mathop{\rm co}C$ for any set $C$ in a Euclidean space; see [36, Theorem 6.3]. ∎
Note that $H$ maps ++$\mathbbm{R}^{\bar{T}}_{++}$ to
@@
i.e. if the conditional support of the day-ahead forward curve contains every positive forward curve. This condition is far from necessary but it could be enforced e.g. by multiplying each $F_{t}(s)$ by $e^{\epsilon_{t}}$, where $\epsilon_{t}$ are independent Gaussian variables with nonzero variance. Such additional randomness could be interpreted as observation errors, as is common in the theory of affine term structure models; see e.g. [10, Section 4]. Taking the variance of $\epsilon_{t}$ sufficiently close to zero, the additional randomness would not have a material effect on portfolio optimization or indifference pricing in the face of bid-ask spreads. More importantly, one cannot trade perfectly liquid zero-coupon bonds with arbitrary maturities in practice, so the no-arbitrage conditions derived above aren’t all that relevant in real-life applications.
@@ 6.2 Arbitrage in futures market
By [34, Theorem 2.43.4], continuity of the exponential function implies
@@
The claim now follows from the fact that $\mathop{\rm ri}\mathop{\rm co}\mathop{\rm cl}C=\mathop{\rm ri}\mathop{\rm co}C$ for any set $C$ in a Euclidean space; see [36, Theorem 6.3]. ∎

### v2

@@ 4.5 Incorporation of user’s views
We apply the calibration procedure from Section 4.2 of [aap21], which allows us to incorporate both short term forecasts as well as long-term views in both the macroeconomic model as well as the forward curve model described in Sections 4.3 and 4.4 above. The calibration is achieved by choosing the sequence of vectors $a_{t}$ in (17) so that the values of $x_{t}$ have desired median values at given points in time. Since both the macro model and the forward curve model are stationary, the model calibration according to Section 4.2 of [aap21] is particularly easy.
@@
Table 11: Long-term median for the macroeconomic risk factors $y$. The $2.5\%$ for $L$ is user-specified. The $2.00\%$ for II corresponds to the inflation target set by the FOMC. The long-term median for $G$ is the median computed from historical data.
$L$ II $G$
$c$ 2.502.50% 2.002.00% 2.852.85%
@@ 5.1 Simulation of the macroeconomy and the forward curve
Figure 8: Simulated scenarios for $L$ (in decimals), II (in percentage), and $G$ (in percentage). The plots show the monthly historical data with a single simulated scenario, the median, and the 95% confidence band.
@@ 5.2 Simulating SOFR derivatives payouts
Table 13: Payouts in USD for long positions in SOFR derivatives. The payouts follow the CME convention where the quantities in the formulas appear to have flipped signs; see Remark 1. The maturities are $t_{1}$ for the futures, $t<t_{1}$ for the call and put, while the maturity of the swaption is denoted by $T$. The strikes are denoted by $X$.
@@
Three-month SOFR futures $F_{t}^{b}(t_{0},t_{1})-R(t_{0},t_{1})]\times 250000$
@@
Put on three-month SOFR futures $P_{t}(t_{2})(F_{t}(t_{1},t_{2})-X)^{+}\times 250000$
@@
If exercised, a swaption, with strike $X$ and maturity $T$, converts to a short position in an OIS with swap rate $X$ and $K$ yearly payments; see [CMEIRprod, CMEswaption]. Assuming perfectly liquid zero-coupon bonds, the market price of such an OIS would be
@@
SOFR futures options deliver the underlying futures contract when exercised. If the underlying futures market is perfectly liquid, we can assume that the options are cash-settled. Indeed, taking the opposite futures position at the time of exercise would yield a cash payout of
$P_{t}(t_{2})(X-F_{t}(t_{1},t_{2}))^{+}\times 250000,$
where $X$ is the option strike and $F_{t}(t_{1},t_{2})$ is the futures price at the time $t$ of exercise.
@@ 5.3 Indifference pricing of SOFR derivatives
A common approach to derivative pricing is to construct a market model together with a martingale measure which is then used for pricing derivatives by taking expectations of their cashflows. Fixing a pricing measure, however, leads to a linear pricing rule, which contradicts the nonlinearities observed in practice. Such problems are addressed explicitly by the techniques of incomplete markets, where one first develops a term structure model under the “statistical” measure and then defines prices according to the indifference principle; see e.g. [car9], [pen14] and their references. This section uses the model developed in Sections 2,  3, and 4 in indifference pricing of the derivatives described in the previous section. The exercise below is quite simple in that the only “hedging instrument” is the money market account represented by the overnight Repo market. More sophisticated hedging strategies will be built in the follow-up paper that takes the stochastic model developed in this paper as the description of the underlying risk factors.
@@
Assuming that the futures options are only exercised at maturity, their indifference prices are given by the same formula.
The indifference price $\alpha$ in (19) can be computed by numerically evaluating the expectations. In the numerical computations below, we will use Monte Carlo with antithetic sampling. That is, we simulate a random sample of $N$ scenarios and obtain another $N$ scenarios by reflecting the underlying Gaussian process.
@@ 6 Arbitrage
The following theorem is from [js98, Theorem 3], but one should note that part g of the statement there has a typo: the symbol $Q_{n}$ should be replaced by $\bar{Q}_{n}$. An alternative proof and an extended statement can be found in [pp24, Remark 2.120]. The ${\cal F}_{t}$}-conditional support of a random variable $X\in L^{0}(\Omega,{\cal F},P;\mathbbm{R}^{d})$ is the smallest closed-valued ${\cal F}_{t}$-measurable mapping $\mathop{\rm supp}\nolimits_{t}X:\Omega\rightrightarrows\mathbbm{R}^{d}$ such that $X\in\mathop{\rm supp}\nolimits_{t}X$ almost surely. The convex hull and the relative interior of a set $C\subseteq\mathbbm{R}^{d}$ will be denoted by $\mathop{\rm co}C$ and $\mathop{\rm ri}C$, respectively.
@@ 6.1 Arbitrage in zero-coupon bonds
By [pp24, Theorem 2.43.4], the continuity of $H$ implies
@@
The claim now follows from the fact that $\mathop{\rm ri}\mathop{\rm co}\mathop{\rm cl}C=\mathop{\rm ri}\mathop{\rm co}C$ for any set $C$ in a Euclidean space; see [roc70a, Theorem 6.3]. ∎
Note that $H$ maps $\mathbbm{R}^{\bar{T}}_{++}$ to
@@
i.e. if the conditional support of the day-ahead forward curve contains every positive forward curve. This condition is far from necessary but it could be enforced e.g. by multiplying each $F_{t}(s)$ by $e^{\epsilon_{t}}$, where $\epsilon_{t}$ are independent Gaussian variables with nonzero variance. Such additional randomness could be interpreted as observation errors, as is common in the theory of affine term structure models; see e.g. [cheridito2007market, Section 4]. Taking the variance of $\epsilon_{t}$ sufficiently close to zero, the additional randomness would not have a material effect on portfolio optimization or indifference pricing in the face of bid-ask spreads. More importantly, one cannot trade perfectly liquid zero-coupon bonds with arbitrary maturities in practice, so the no-arbitrage conditions derived above aren’t all that relevant in real-life applications.
@@ 6.2 Arbitrage in futures market
By [pp24, Theorem 2.43.4], continuity of the exponential function implies
@@
The claim now follows from the fact that $\mathop{\rm ri}\mathop{\rm co}\mathop{\rm cl}C=\mathop{\rm ri}\mathop{\rm co}C$ for any set $C$ in a Euclidean space; see [roc70a, Theorem 6.3]. ∎